OpenStudy (anonymous):

if the sales tax is 3% what is the cost of a pair of shoes with a price of $55 ILL GIVE U A MEDAL PLZ

OpenStudy (anonymous):

total price = cost + sales tax

OpenStudy (anonymous):

ANSWER ?

OpenStudy (anonymous):

you have to go for the technique not for the answer

OpenStudy (anonymous):

IS IT 58 ?

OpenStudy (anonymous):

how you got 58. if you show me your work then i can tell you whether it is correct or not

OpenStudy (anonymous):

I divide

OpenStudy (anonymous):

divide what by what

OpenStudy (anonymous):

wait no I just add 3

OpenStudy (anonymous):

165

OpenStudy (anonymous):

220

OpenStudy (anonymous):

ok let me do it for you

OpenStudy (anonymous):

lol alrightt

OpenStudy (anonymous):

that's not one of my answer choices :(

OpenStudy (anonymous):

ok i got my mistake

OpenStudy (anonymous):

what Is it?

OpenStudy (anonymous):

let the price of the the shoes is x then total price = cost + sales tax 55 = x +0.03x 55 = x(1+0.03) 55 = 1.03x x = 55/1.03 = 53.39 cost = 53.39$

OpenStudy (anonymous):

nope the closest one to that is 56.65

OpenStudy (anonymous):

is the question you posted is correct???? please check it agin

OpenStudy (anonymous):

yessir

OpenStudy (anonymous):

basically the cost should be less than the price of the shoes other wise it make no sense when the cost become larger than the price

OpenStudy (anonymous):

shouldn't it be more since tax is adding ! like wen u buy something doesn't tax add

OpenStudy (anonymous):

yes but tax increases the price no the cost

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

isn't price and cost the same thing

OpenStudy (anonymous):

or basically in the question it not asked to find the cost it is asking about to find the price while the cost is 55

OpenStudy (anonymous):

no no they are completly the different things

OpenStudy (anonymous):

ughhh alright

OpenStudy (anonymous):

if it is the question then 56.65 is your right answer

OpenStudy (anonymous):

alright thank u

OpenStudy (anonymous):

you are welcome but please make sure to post the questions correctly next time

OpenStudy (texaschic101):

.03 x 55 = 1.65 (this is your sales tax) 55 + 1.65 = 56.65 (total cost of shoes including sales tax)